### v1

# https://orcid.org/0009-0005-6083-0932

Abstract
Recent studies have explored the performance of Large Language Models (LLMs) on various Software Engineering (SE) tasks, such as code generation and bug fixing. However, these approaches typically rely on the context data from the current snapshot of the project, overlooking the potential of rich historical data from real-world software repositories. Additionally, the impact of prompt styles on LLM performance within a historical context remains underexplored. To address these gaps, we propose HAFix, which stands for History-Augmented LLMs on Bug Fixing, a novel approach that leverages individual historical heuristics associated with bugs and aggregates the results of these heuristics (HAFix-Agg) to enhance LLMs’ bug-fixing capabilities. To empirically evaluate HAFix, we employ Code Llama on a dataset of 51 single-line bugs, sourced from 11 open-source projects, by mining the historical context data of bugs and operationalizing this context in the form of seven heuristics. Our evaluation demonstrates that historical heuristics significantly enhance bug-fixing performance. For example, the FLN-all heuristic achieves a 10% improvement in performance compared to a non-historical baseline inspired by GitHub Copilot. Furthermore, HAFix-Agg fixes 45% more bugs than the baseline, outperforming FLN-all and demonstrating the best performance overall. Moreover, within the context of historical heuristics, we identify the Instruction style prompt as the most effective template for LLMs in bug fixing. Finally, we provide a pragmatic trade-off analysis of bug-fixing performance, cost, and time efficiency, offering valuable insights for the practical deployment of our approach in real-world scenarios.
Keywords:
Bug fixing Large Language Model Software development history LLM cost analysis
1 Introduction
Large language models (LLMs) have emerged as transformative tools in software engineering (SE), with applications spanning code generation and completion (Lu et al., 2021; Chen et al., 2021; Du et al., 2023; Yu et al., 2024; Li et al., 2024b; Zhang et al., 2024a), and bug fixing (Jiang et al., 2023; Zhang et al., 2024b; Fan et al., 2023; Xia et al., 2023; Li et al., 2024a; Jimenez et al., 2023; Aleithan et al., 2024). These models, such as CodeBERT (Feng et al., 2020), CodeT5 (Wang et al., 2021), Codex (Chen et al., 2021), and Code Llama (Roziere et al., 2023), have demonstrated significant capabilities in understanding and generating code based on various contextual inputs. The main key to unlocking the potential of LLMs is finding ways to more effectively leverage the full spectrum of context available in software development, particularly when it comes to understanding and resolving bugs.
While recent advances have focused on evaluating LLM performance by utilizing contextual information such as buggy-line-surrounded function code snippets to guide LLMs in bug fixing (Ahmad et al., 2021; Wang et al., 2021; Niu et al., 2022; Xia and Zhang, 2022; Jiang et al., 2023; Fan et al., 2023; Xia et al., 2023; Jimenez et al., 2023; Xia and Zhang, 2024; Zirak and Hemmati, 2024), the potential of incorporating historical context remains underexplored, except for Le et al. (2016)’s use of previously-appearing fix patterns from different projects to guide the current bug fixing. Historical data, such as information from blame commits, encapsulates the incremental evolution of a project, reflecting the developers’ intent, bugs origins, and the rationale behind previous fixes. This historical information draws inspiration from early MSR (Mining Software Repositories) works, which emphasized the importance of historical data in understanding the evolution of software bugs (Śliwerski et al., 2005; Hassan, 2006). These insights can be instrumental in understanding the context of bugs and guiding LLM in bug-fixing strategies. However, there is still a significant gap in leveraging rich history data to help LLMs with bug fixing.
Moreover, prompt design is another critical factor influencing LLM performance in bug fixing, as the way in which information is presented to LLMs can significantly influence the relevance and quality of the generated code (Sclar et al., 2023). While recent work (Jiang et al., 2023) studied if the LLMs can make good use of the buggy line, they did not systematically evaluate how different prompt styles perform when incorporating the historical context of a bug alongside natural language instructions. This highlights the need to explore which prompt style works best in leveraging historical heuristics to enhance LLM bug-fixing performance.
While historical heuristics may enhance bug-fixing performance, they also come with increased prompt size, monetary cost, and longer inference time. This raises the need to explore how historical data can be effectively utilized in a cost-efficient manner with LLMs. Although prior work, such as Xia and Zhang (2024), has examined the price cost of ChatGPT for fixing a single bug, it did not consider the role of historical data in LLM-based bug fixing, i.e., the increased prompt because of the rich historical context information. Similarly, Jiang et al. (2023) provided the analysis of the model size and its relationship with performance, without delving into the broader implications of incorporating historical heuristics for LLMs on bug fixing. In other words, a comprehensive exploration of the trade-offs between bug-fixing performance, inference price, and time efficiency when leveraging historical heuristics remains missing.
Based on these research gaps, we aim to investigate how historical context, particularly blame commit data, can enhance LLM performance in bug fixing. Additionally, we aim to evaluate three distinct prompt styles: Instruction (Instruction), Instruction with the buggy line labeled (InstructionLabel), and Infill (InstructionMask), to assess their impact on bug-fixing effectiveness when used with historical context. Furthermore, our objective is to analyze the trade-offs between computational cost, time efficiency, and bug-fixing performance to provide actionable insights into the practical use of historical data in LLM-based bug fixing.
Based on these objectives, we determine the following research questions to explore:
  • RQ1: Do history-augmented LLMs improve bug fixing compared to models without historical context?
  • RQ2: How do different prompt styles impact the bug-fixing performance of history-augmented LLMs?
  • RQ3: What is the cost and efficiency of history-augmented LLMs on bug fixing?
To address our research questions, we propose HAFix, which stands for History-Augmented LLMs on Bug Fixing, a novel approach that leverages individual historical heuristics associated with bugs and aggregates the results of these heuristics (HAFix-Agg) to enhance LLMs’ bug-fixing capabilities. To empirically evaluate HAFix, we mined and constructed a dataset of 51 single-line Python bugs from BugsInPy (Widyasari et al., 2020), focusing on real-world projects with rich development histories. For RQ1, we explored the integration of historical data by mining and categorizing historical data from blame commit and designing different historical heuristics to augment our baseline. The baseline was inspired by how GitHub Copilot (Copilot, 2024b) constructs prompts in practice (Copilot, 2024a). We employed a Code Llama 7b instruct-tuned model as our subject model. The pass@k metric was used to assess the likelihood of generating correct results across multiple model outputs.
For RQ2, we investigated the influence of prompt styles on bug-fixing effectiveness. Using the baseline and the top-performing approaches from RQ1 (FLN-all and HAFix-Agg), we systematically tested three distinct prompt styles: Instruction, InstructionLabel, and InstructionMask. The results identified the optimal combination of prompt style and historical context.
For RQ3, we analyzed the cost and efficiency of inference using HAFix-Agg with the Instruction prompt style, as it was the best-performing configuration from RQ2. We measured inference time by recording the duration required to generate 10 outputs via nucleus sampling, ensuring consistency by running all experiments on identical infrastructure. To estimate the inference price, we used a cost model derived from GPT-3.5-turbo-instruct’s pricing (OpenAI, 2024), factoring in token counts for input and output. Additionally, we defined four execution scenarios: Exhaustive, EarlyStop (ES), ES-AccSorted, and ES-UniSorted, to explore the trade-offs between bug-fixing performance, computational cost, and time efficiency.
This study makes the following contributions:
  • Leveraging Historical Context for LLM-based Bug Fixing: This work investigates the impact of integrating various history-augmented heuristics derived from blame commits into LLM prompts to evaluate bug-fixing performance. It introduces an innovative approach called HAFix, demonstrating how leveraging historical context enhances bug-fixing performance and expands the model’s capability to resolve complex bugs. Our findings show that one heuristic (FLN-all) achieves a 10% improvement and HAFix-Agg improves bug-fixing performance by 45% compared to the baseline while addressing all bugs fixed by the baseline.
  • Comprehensive Prompt Style Evaluation: We comprehensively analyze three distinct prompt styles including Instruction, InstructionLabel, and InstructionMask, and reveal their respective impacts on LLM bug-fixing performance. This study identifies the Instruction prompt as the most effective style for leveraging historical context.
  • Pragmatic Performance-Cost-Efficiency Analysis: We provide an in-depth evaluation of the trade-offs between bug-fixing performance, inference price, and time efficiency across various historical heuristics and their execution sequences. Strategies such as the early stop of heuristic execution are highlighted for effectively balancing efficiency and effectiveness.
Our contributions lay the groundwork for leveraging historical data and optimal prompt design to improve LLM-based generated code and bug fixing, providing actionable guidance for balancing performance and cost in real-world applications.
2 Related Work
2.1 Usage of LLMs in the Context of Software Engineering
Large Language Models (LLMs) have rapidly become a valuable tool in software engineering (SE), enabling various tasks such as code generation, code completion, and automated program repair (APR). Recent advancements in LLMs, including models like CodeBERT (Feng et al., 2020), CodeT5 (Wang et al., 2021), Codex (Chen et al., 2021), and Code Llama (Roziere et al., 2023), have shown strong performance in understanding and generating code based on local and repository-level contexts.
2.1.1 Code Generation
In the context of code generation, recent works have explored various benchmarks and methods to enhance and evaluate LLM performance. RepoBench (Liu et al., 2023) retrieves the most relevant code snippets from other files for code completion, but all retrieved context information is from the current snapshot of the project. Similarly, ClassEval (Du et al., 2023) focuses on class-level code generation, identifying limitations in how LLMs handle class structures and dependencies. Additionally, CoderEval (Yu et al., 2024) introduces a benchmark for pragmatic code generation, highlighting areas where LLMs need improvement to generate functional, maintainable code for real-world applications. The DevEval (Li et al., 2024b) benchmark assesses models like GPT-4 and Code Llama on real-world software projects, revealing challenges in generating practical code. RepoHyper (Phan et al., 2024) improves code completion by constructing semantic graphs, allowing LLMs to prioritize relevant code snippets but still without considering history context. CodePlan (Bairi et al., 2024) introduces a planning-based approach, where LLMs generate sequences of code edits based on context from the current project snapshot, showing promise in large-scale code modifications.
2.1.2 Automated Bug Fixing
In the LLM-based bug fixing field, most works focus on single-line bugs and providing models with the buggy-line surrounded code snippets (Lu et al., 2021; Guo et al., 2020; Ahmad et al., 2021; Chakraborty and Ray, 2021; Wang et al., 2021; Zhang et al., 2022; Niu et al., 2022; Chakraborty et al., 2022). Recent works have investigated various methods to enhance the effectiveness of LLMs in fixing software bugs. Fan et al. (2023) have shown that given proper instructions such as information from fault localization, LLMs show promising results and can outperform traditional bug-fixing tools. Xia et al. (2023) highlighted the importance of leveraging fine-tuning and prompting to harness the power of LLM with identifiers extracted from lines that are very similar to the buggy line. The SWE-bench benchmark (Jimenez et al., 2023) evaluates LLMs on real-world GitHub issues, identifying that while LLMs can resolve straightforward bugs, they often falter on complex, context-dependent issues. Jiang et al. (2023) demonstrated that while LLMs show promising results, they struggle to effectively utilize the buggy line, but when fine-tuned, they exhibit enhanced bug-fixing capabilities, although they may potentially over-rely on the buggy line. Furthermore, Hossain et al. (2024) conducted a deep dive into bug localization and repair, localizing and fixing bugs at the token granularity rather than the traditional line granularity, resulting in substantial improvements in bug-fixing performance.
More recent works start considering the repository-level context information but only from the current snapshot of the project. RepoBugs (Chen et al., 2024) introduces repository-level benchmarks, revealing that LLMs perform better when provided with extensive repository-level context. In another work, Prenner and Robbes (2024) examine the impact of local context from the current snapshot of the project in neural program repair, revealing that increasing context size significantly improves performance while emphasizing the need for clear context documentation and adequate datasets. Furthermore, domain adaptation techniques have been proposed to align models with specific codebases (Zirak and Hemmati, 2024), enhancing repair success rates, while hybrid approaches (Li et al., 2024a) combining LLMs with program analysis provide promising results in generating more reliable fixes. Finally, Zhang et al. (2024b) highlight the integration of diverse software engineering agents to enhance the effectiveness of LLMs in solving real-world GitHub issues, utilizing the same contextual input across different agents.
One of the state-of-the-art bug-fixing tools in practice is GitHub Copilot (Copilot, 2024b). According to the official prompt engineering guidelines of GitHub Copilot (Copilot, 2024a), when generating code suggestions, it uses the lines immediately before and after the user’s current cursor position, as well as information from other files open in the editor and the URLs or file paths to provide relevant context. This information is derived solely from the current project snapshot, without incorporating historical data.
Unlike previous research that primarily focuses on leveraging context information from the current snapshot of the project for code generation and automated bug fixing, our work explores the largely untapped potential of historical context from previous snapshots of the project in enhancing LLM performance for software engineering tasks. While recent studies have expanded the input context window of LLMs, utilized repository-level information, or employed domain adaptation techniques, they have not incorporated the code’s evolutionary history to inform bug fixing or code generation. Our study addresses this gap by systematically evaluating the impact of historical context on LLM-based automated bug fixing, providing insights that could generalize to other software engineering tasks where the history of code evolution is a crucial factor. To evaluate the impact of historical context, we design a prompt inspired by GitHub Copilot’s practices as a baseline, reflecting traditional bug-fixing methods. The details of the baseline prompt and its implementation are discussed in Subsection 4.2.1.
2.2 LLM Prompt Style vs. Bug Fixing
There is limited research on prompt styles for large language models (LLMs) in bug fixing, yet LLMs demonstrate sensitivity to prompt template choices (Sclar et al., 2023). The work by Xia and Zhang (2022) introduces a cloze-style APR approach that directly leverages LLMs without requiring any fine-tuning or retraining on bug-fix datasets, framing the repair process as a cloze task to predict masked code snippets. Additionally, the study by Jiang et al. (2023) evaluates various code language models (CLMs) for APR, determining whether to explicitly label the buggy line or mask it in the prompt template based on the corresponding pre-training task style of the LLM. Furthermore, Xia and Zhang (2024) present a conversation-driven approach that employs cloze-style prompts, interspersing patch generation with immediate feedback to enhance the interaction between the model and the repair task. Recent work by Sclar et al. (2023) emphasizes the critical nature of prompt formatting, demonstrating that even minor variation can lead to drastic performance changes.
In contrast to these studies, our research evaluates both the usage of instruction and infill prompts, to measure the impact of various prompt styles (instruction, instruction with buggy lines labeled, and infilling) on LLM performance in bug fixing. Our analysis focuses on how these prompt styles affect the model’s effectiveness when applied in conjunction with various history heuristics. This investigation offers a unique perspective on the relationship between prompt design and bug repair performance, contributing to a more nuanced understanding of how customized prompts can enhance LLM capabilities in bug-fixing.
2.3 LLM Cost vs. Bug Fixing
Recent research has increasingly focused on analyzing the cost implications of using large language models (LLMs) for bug fixing. Jiang et al. (2023)’s study on the impact of code language models on APR examines the trade-off between model size and bug-fixing capability, showing that while larger models offer higher success rates, they also incur greater computational costs. Additionally, Xia and Zhang (2024) introduce a conversation-driven APR approach using ChatGPT, achieving an average repair cost of \$0.42 per bug, emphasizing the cost-effectiveness of using conversational LLMs in automated bug fixing. Hidvégi et al. (2024) propose a cost-efficient program repair method that minimizes token costs by optimizing prompts and leveraging strategies like summarizing responses and patch multiplication while maintaining high bug-fixing performance. Similarly, Nayab et al. (2024) explore how the length of LLM-generated outputs influences both inference cost and model performance, offering strategies to minimize unnecessary token generation for more cost-effective results. Shekhar et al. (2024) optimize LLM usage costs by predicting output quality and selecting models to balance quality, cost, and latency, showing significant improvements in cost-efficiency and quality.
In contrast, our work goes beyond just analyzing inference price by also examining inference time and their trade-offs with bug-fixing performance. Unlike prior studies that focus solely on cost or performance in isolation, we investigate how our history-augmented bug-fixing approach balances high performance with both cost and time efficiency. Furthermore, we investigate the cost of using different historical heuristics for LLM-based bug fixing, providing a pragmatic cost estimation by referencing GPT-3.5-turbo-instruct’s (OpenAI, 2024) pricing model. This holistic approach addresses a gap in the current literature, offering insights into the large-scale, practical application of LLMs in software engineering.
3 HAFix: History-Augmented LLMs for Bug Fixing
We introduce HAFix (History-Augmented LLMs on Bug Fixing), a novel approach that enhances bug-fixing capabilities by incorporating historical heuristics extracted from blame commit data. By integrating historical data into the prompts, HAFix provides the model with additional context, aiding in identifying the root cause of the bug and generating a possible solution to solve it. Yet, what is the most relevant historical context data for a bug? We explore this question from the perspectives of temporal and spatial analysis.
The spatial aspect of bug fixing involves understanding the structural and positional context within the codebase. This approach, commonly utilized by tools like GitHub Copilot (Copilot, 2024b), emphasizes using information from the current snapshot of the code, such as the surrounding function code and file structure (Copilot, 2024a). By focusing on the specific function and narrowing it down to the buggy line, the model is provided with the most relevant spatial context, minimizing irrelevant data. This approach prevents the model from being distracted by unrelated sections of the code and ensures that the model’s attention is directed at the precise location of the bug. For example, file names and function-level code surrounding the buggy line provide context to pinpoint where the bug exists and how it might be fixed.
From the temporal perspective of a bug, the commit that last touches the buggy line (blame commit) will give the most closely related information about how this buggy code is modified (Śliwerski et al., 2005; Hassan, 2006). This information includes details of the changes made, the reasoning behind these changes, and the broader context of other modifications within the same commit. This temporal analysis draws inspiration from early MSR (Mining Software Repositories) techniques, which emphasized the importance of historical data in understanding the evolution of software bugs. By integrating both spatial and temporal data, HAFix combines established practices with novel insights to improve bug-fixing capability.
Figure 1: Dataset collection for HAFix: ① represents the data used for the baseline, while ② to ⑧ represent the data for various historical heuristics. V4 refers to the snapshot of the project version where the bug fix was committed, and V3 is the snapshot of the previous version containing the bug. V2 is the snapshot of the last commit modifying the buggy line in the V4 snapshot, while V1 is the snapshot of the commit preceding V2. The rationale for selecting the blame commit and these historical heuristics are detailed in Section 3.
Figure 1 illustrates the data collection process for HAFix, which is structured across multiple stages corresponding to different project versions (V1, V2, V3, V4). The data collection process begins at V4, which is the snapshot of the project version where the fixed code was committed (fix commit). Next, we trace back to V3, which is the version preceding V4 that still contains the buggy code (buggy commit). Using PyDriller (Spadini et al., 2018), we then identify V2 as the version that last modifies the buggy code line (blame commit). Lastly, we trace back to V1, which is the snapshot right before V2, to pair it with V2 as the historical data of the blame commit. V4 and V3 capture the direct and actual changes made to resolve the bug, which will be used for the baseline detailed in the following Subsection 4.2.1. V2 and V1 contain the historical data of the bug, we will use these different heuristics to enhance the baseline, which will be detailed in Subsection 3.1.
3.1 History-Augmented Bug Fixing: Historical Heuristics Prompt
The key innovation in HAFix lies in the augmentation of bug-fixing prompts with historical context. We extract this context by mining the blame commit, which represents the last modification of the buggy line. This process provides temporal insights that highlight how the code evolved, shedding light on potential root causes of the bug. As shown in Figure 1, to construct history-augmented prompts, we extract the following seven heuristics from the blame commit (V2) and the previous commit (V1) including:
  • Co-evolved Functions’ Names in the Modified Buggy File (CFN-modified): The names of functions modified within the buggy file in commit V2. These are important because they provide context about which specific functions were altered and may directly influence the buggy line.
  • Co-evolved Functions’ Names in All Modified Files (CFN-all): The names of functions modified across all modified files in commit V2. This information helps in understanding broader structural changes within the codebase that may indirectly impact the buggy line.
  • All Functions’ Names in the Modified Buggy File (FN-modified): The names of all functions (whether changed or not) in the modified buggy file in the commit of V2. This allows the model to understand the function structure in the file, offering context to better locate the bug within its function.
  • All Functions’ Names in All Modified Files (FN-all): The names of all functions (whether changed or not) in all modified files in the commit of V2. This information helps capture a wider scope of the code changes and potential interactions between different functions across the codebase.
  • Co-evolved Files’ Names (FLN-all): The names of changed files of the commit in V2. This provides a broad context of which files were modified, potentially indicating areas in the code that might affect the buggy line or the system’s behavior.
  • Function Code Pairs (FN-pair): The function code before and after the blame commit in V1 and V2. This temporal context helps identify how the buggy function evolved and provides clues about what changes directly contributed to the bug’s introduction.
  • File Diff Patch (FL-diff): The diff patch from the git diff command in the commit of V2. This allows us to see the exact code changes made, providing precise details on what was modified, which can aid in pinpointing the cause of the bug.
These heuristics were chosen to offer a comprehensive yet focused historical snapshot, essential for understanding both the bug’s cause and the code structure surrounding it.
3.2 HAFix-Agg: Aggregated HAFix Variant
Embedding all heuristic data directly into a single prompt would result in excessively large inputs, increasing both computational cost and latency. To address this, we introduce HAFix-Agg, a variant that aggregates the LLM results of individual heuristics, which helps reduce the prompt size and computational cost of model inference. While HAFix-Agg aims to improve performance by leveraging insights from multiple heuristics, it comes at the expense of higher inference costs, as each heuristic requires separate inference runs. This trade-off allows for broader coverage of potential fixes, making HAFix-Agg particularly useful for complex bug-fixing scenarios where a single heuristic may be insufficient. We assess this variant’s prediction performance (RQ1/2) and further explore its cost-effectiveness in RQ3, providing insights into its feasibility and scalability for real-world deployment.
4 Empirical Evaluation of HAFix
To empirically validate the effectiveness of HAFix, we conduct a comprehensive evaluation using real-world Python bugs. This section details our dataset selection, data collection process, model selection, prompt construction, experimental pipeline, and inference infrastructure.
4.1 Dataset Selection
In this study, we focus on bug datasets consisting of Python source code. This decision stems from the observation that most current LLM evaluation studies on bug fixing center around Java bugs (Lu et al., 2021; Xia and Zhang, 2022; Jiang et al., 2023; Zirak and Hemmati, 2024), despite Python being one of the most popular programming languages, even surpassing Java in popularity (Jaydevs, 2023). Hence, we aim to contribute to this field by helping to narrow the gap in bug fixing for Python compared to Java. Additionally, we require a dataset that includes real-world projects with test cases, allowing us to verify the functional correctness of model-generated code against the developer’s original fixed code. Lastly, we prioritize selecting datasets that contain popular projects with rich development history, enabling us to mine the historical context data for each bug.
We select BugsInPy collected by Widyasari et al. (2020) as our subject benchmark dataset for several reasons. BugsInPy is a comprehensive, hand-curated dataset with 493 real-world bugs from 17 large, non-trivial Python projects. The bugs in BugsInPy are carefully selected to meet specific criteria: they must involve changes in the source code, excluding modifications like configurations or build scripts. Additionally, the bugs should be reproducible, with at least one test case failing on the faulty version, and they must be isolated from unrelated changes, such as refactoring or feature additions. This ensures the quality of the bugs for our study. The projects included in BugsInPy span various domains such as machine learning, developer tools, scientific computing, and web frameworks. This diversity is crucial as it allows us to do our evaluation across various real-world scenarios, making our findings more generalizable.
Table 1: Summary of subject bugs and their associated project source information as of November 2024.
Project Name Number of Bugs Lines of Code Number of Stars
sanic [^1] https://github.com/huge-success/sanic.git 1 77k 18.1k
luigi [^2] https://github.com/spotify/luigi.git 9 44k 17.9k
youtube-dl [^3] https://github.com/ytdl-org/youtube-dl.git 7 139k 132k
ansible [^4] https://github.com/ansible/ansible.git 2 237k 63k
scrapy [^5] https://github.com/scrapy/scrapy.git 4 479k 53.1k
pandas [^6] https://github.com/pandas-dev/pandas.git 15 457k 43.8k
thefuck [^7] https://github.com/nvbn/thefuck.git 6 11k 85.4k
tornado [^8] https://github.com/tornadoweb/tornado.git 2 29k 21.7k
fastapi [^9] https://github.com/tiangolo/fastapi.git 1 165k 77.6k
black [^10] https://github.com/psf/black.git 1 118k 39.2k
tqdm [^11] https://github.com/tqdm/tqdm.git 3 7k 28.7k
Following previous work (Xia and Zhang, 2022; Ye et al., 2022; Jiang et al., 2023; Prenner and Robbes, 2024), we also focus on single-line bugs from BugsInPy, i.e., bugs whose fixes are focused on one line of code. To the best of our knowledge, limited research has specifically evaluated LLM’s performance in fixing single-line Python bugs. Furthermore, by starting with single-line bugs, we can establish a solid foundation before tackling more complex cases, such as multi-line or even multi-hunk bugs (i.e., bugs where the faulty code lines are not contiguous) in future work.
To identify single-line bugs in BugsInPy, we examined the code changes in the commit that fixed the bug (the “fix commit”). Specifically, we used the open-source tool PyDriller (Spadini et al., 2018) along with the fix commit ID provided by BugsInPy to locate the fix commit. We then verified whether the commit contained only one change in a single Python file, excluding test files. Further, we checked if the intersection of added and deleted lines involved a single line of code change while excluding no-code lines such as blank lines or comments. Through this process, we identified 68 single-line bugs from a total of 493 diverse and representative bugs.
We manually validated each single-line bug to ensure it met the specified criteria above. With the location and isolation of bugs confirmed during the initial identification above, our primary focus was verifying reproducibility. This step involved running test cases to confirm they can pass in the fix commit and fail in the buggy commit (the immediate predecessor). For example, if the test cases of a bug pass in both the fixed and buggy commits or if they fail in both, we filter out such cases.
Ultimately, out of the original 68 bugs, we obtained a subject dataset of 51 high-quality single-line bugs, which is a similar dataset size as prior works (Prenner et al., 2022; Kolak et al., 2022; Fan et al., 2023; Peng et al., 2024; Chen et al., 2024). Table 1 lists the project sources of these 51 bugs, including their number of lines of code and GitHub star counts. We believe that the popularity of these projects highlights the representativeness of these bugs. For each bug, we rely on the fix commit ID from BugsInPy to locate the corrected code and its corresponding test cases. We also use this commit ID to trace back to the buggy commit (the immediate predecessor) and mine the necessary data for our study, which will be detailed below.
To provide a deeper understanding of the dataset and demonstrate the effort involved in its curation, we selected one representative example from the 51-bug dataset. This example was chosen to highlight the diversity and complexity of the bugs included in the dataset. We provide detailed information in the example, including the commit description, heuristic values, and relevant metadata. Full details of the selected example can be found in the Appendix B.
4.2 Data Collection
The data collection process follows the staged approach depicted in Figure 1, progressing from the most recent fix commit (V4) to the earliest commit (V1).
4.2.1 Baseline Data Collection
To establish a baseline for comparison, we need to design a prompt that reflects traditional bug-fixing practices without incorporating historical data. As mentioned in Subsection 2.1.2, this baseline prompt serves as a reference point for evaluating the effectiveness of history-augmented approaches. We design our baseline prompt inspired by how GitHub Copilot (Copilot, 2024b) processes user prompts, as it is one of the most widely used coding assistant tools for bug fixing in practice. However, the specific prompt template used by GitHub Copilot is not publicly available. According to the official prompt engineering guidelines of GitHub Copilot (Copilot, 2024a), when generating code suggestions, it uses the lines immediately before and after the user’s current cursor position, as well as information from other files open in the editor and the URLs or file paths to provide relevant context. This information is derived solely from the current project snapshot, without incorporating historical data.
For each bug in our subject dataset, we mined the non-history data for constructing the baseline prompt from the fix commit (V4) and buggy commit (V3), providing spatial context such as function-level code and the buggy line without temporal information. The fields we mined include:
  • Project Name: This field provides the LLM with the repository name associated with the bug.
  • Buggy File Name and Path: Specifies the name and path of the buggy file that was modified to fix the bug, crucial for locating the bug within the project’s codebase. This field remains consistent between V4 and V3, as no file renamings were observed in the studied bugs.
  • Buggy Line Location: Since our focus is on bug repair rather than fault localization, we explicitly provide the LLM with the precise buggy line code. The buggy line code should remain consistent across all commit snapshots from V2 to V3, we extract it only from V3 for simplicity.
  • Buggy Function Name: Indicates the specific function where the bug was located, providing more precise localization within the file. The field should be the same in V4 and V3.
  • Function Code Before and After the Fix Commit: It provides LLM with the whole buggy line surrounded with the function-level code before and after the fix, allowing for a detailed examination of the changes at the function level. The function code before the fix commit is from V3 and after is from V4.
  • Bug Description: This field provides the essential bug context, which will be detailed in Subsection 4.2, using the cleaned-up version mined from GitHub issue pages or commit messages, ensuring no post-fix details were included to prevent data leakage.
We also provide an example of the baseline prompt built based on data collected above in Listing 2 in Appendix A.
Note that we do not mine the entire buggy or fixed file code, but instead narrow the scope to the function-level code snippet surrounding the buggy line, since this provides sufficient context to understand the single-line bug while avoiding noise from unrelated parts of the file. Additionally, we employ AST (Abstract Syntax Tree) matching for precise localization of the function-level code snippet, since AST offers a structural representation of the source code, enabling us to identify the exact code elements that changed. For instance, in cases where multiple functions share the same name within a file, AST matching distinguishes these functions based on their unique parameter combinations and parent nodes, such as a file or class. Moreover, for the bug description field, we consider mining both the commit message and the corresponding GitHub issue page as detailed below.
Bug Description Mining
In practice, before fixing a bug, developers often have access to contextual information such as how the bug manifests, its consequences, and any error output. Therefore, it is crucial to incorporate this context into our approach to ensure a comprehensive evaluation. However, fix commit messages typically lack this level of detail. To supplement this, we mined the corresponding GitHub issue page for each commit. We began by manually identifying whether the commit message contained a link to the relevant GitHub issue page; when such a link was present, we recorded it for the bug. Only if no link was present, we used the commit message to derive the bug description for each commit, as a GitHub issue page generally provides detailed descriptions, discussions, and the steps taken to resolve the bug. Using the open-source tool GHApi (ghapi, 2024), we extracted the title and body (the initial comment block) of the issue page, which provides a concise yet informative bug description. Figure 2 provides an example of the bug description information that we mined from the GitHub issue page. We combine the issue title and body, highlighted within the two red-circled boxes, to create the bug description.
Figure 2: An example of the bug description we mined from the GitHub issue page.
To avoid data leakage in our empirical evaluation, we have to ensure that the bug description that we mined from GitHub issue pages or fix commit messages does not contain post-hoc details about how a bug was fixed. To avoid such an issue, we manually checked each bug description and filtered the information that was too closely related to the fix for the bug. Our goal was to ensure that the model focused on understanding the general context and nature of the bug, without exposure to the exact fix, which could otherwise compromise our evaluation of the model’s ability to independently generate a correct solution.
4.2.2 Historical Data Collection (HAFix Heuristics)
While non-history data provides valuable spatial context, it lacks insights into the evolution of the bug. To bridge this gap, we incorporate historical data collection to enhance the prompts with temporal context. For each bug in our subject dataset, we mined the history data for constructing the prompt of each HAFix heuristics from the blame commit (V2) and previous commit (V1), providing buggy line-related temporal context such as co-evolved files, functions, and diffs. The fields we mined follow the Subsection 3.1.
The combination of non-history and historical data collection forms the foundation of HAFix, enriching bug-fixing prompts with both spatial and temporal perspectives to improve LLM performance in bug fixing.
4.3 Model Selection
For our experiments, we selected the 7b instruct-tuned version from the Code Llama family of models. The choice of this model was influenced by several key factors that align with our project’s requirements and constraints.
One of the primary considerations was the model size. The 7b instruct-tuned variant has a manageable size of 12.55GB, which strikes a balance between performance and resource efficiency. In contrast, the larger 13b and 34b variants, with sizes of 24GB and 63GB respectively, require significantly more computational resources and memory. For our experimental setup, the 7b variant was deemed more practical. It allows us to run the model efficiently without compromising too much on the quality of the generated code.
Another critical factor in our model selection was the instruct-tuning and code-infilling capabilities. The Code Llama 7B Instruct model is specifically designed to understand and follow instructions, making it particularly suitable for generating code based on the structured prompts we provide. Additionally, its ability to perform code infilling allows it to complete code segments given surrounding context, enhancing its utility in our tasks.
While the Code Llama family also includes a base variant and a Python-tuned variant, neither was as appropriate for our needs. The base variant, lacking instruction tuning, may not interpret and respond to our prompts as effectively as the instruct-tuned version. The Python-tuned variant, while optimized for Python code generation, is not specifically tuned for understanding and executing instructions, which could lead to suboptimal performance in our instruction-driven tasks.
4.4 Prompt Construction
Figure 3: Example of three prompt styles.
We designed and evaluated three different prompt styles to provide input to the LLM to fix bugs, as different prompting styles can influence the outputs generated by LLMs (Sclar et al., 2023). Each prompt style is applicable to the baseline and each heuristic. Figure 3 provides an example for three prompt styles. These styles are categorized as follows:
  • Instruction: This prompt style presents the entire buggy code snippet and highlights the buggy line in the instruction text. It includes the function with the bug, along with the full implementation context. In Figure 3, the first box demonstrates this style, with the buggy line highlighted in the instruction text at the bottom.
  • InstructionLabel: This prompt style labels the buggy line within the function code to provide more precise guidance. The buggy line is tagged with <BUGGY_LINE>, directing the LLM’s focus to the specific part of the code that needs fixing. The second box in Figure 3 showcases this style, where the buggy line is labeled and tagged in both the function and the instruction.
  • InstructionMask: This prompt style masks the buggy line with a placeholder, <FILL_ME>, and highlights it in the instruction text. The LLM then generates the correct code to replace the masked line. The third box in Figure 3 illustrates this style, with the masked line in the function and the corresponding instruction guiding the model to generate a fix.
The rationale for using the Instruction is that previous LLM-based bug-fixing approaches only provide the model with buggy code (Lu et al., 2021). However, most LLMs, including our subject model, Code Llama 7b instruct, are trained to understand both natural and programming languages. For InstructionLabel, we were inspired by previous work (Jiang et al., 2023) that explicitly labeled the buggy line within the functional code. Finally, we evaluated InstructionMask based on Code Llama’s capacity for infilling tasks and its original design for code completion. However, to the best of our knowledge, a comparison between prompt styles has not been conducted on Code Llama. Inspired by this, we explored whether masking and regenerating the buggy line could yield better performance than direct fixes.
4.5 Experimental Pipeline
With the selected model and prepared prompt styles, we can now feed the prompts to the model and initiate the experimental pipeline. Figure 4 shows an overview of our experiment pipeline, which we describe step by step.
Figure 4: Experiment pipeline for HAFix evaluation.
  1. 1.
    Constructing and Categorizing Prompts. The first step in our experimental pipeline is to construct prompts in three distinct styles. As illustrated in Figure 4, the Instruction prompt group includes a baseline (see Subsection 4.2.1) and seven history-based HAFix heuristics, each variant enriched with varying historical information mined from bug blame commit data (see Subsection 4.2.2). In both the baseline and these historical heuristics variants, the buggy code is presented in the Instruction style. Additionally, for InstructionLabel and InstructionMask styles, the buggy line and code are presented in the baseline and heuristics by following the corresponding structure as described in Subsection 4.4.
  2. 2.
    Feeding Prompts to Code Llama. The constructed prompts are then input into the Code Llama model (Roziere et al., 2023) to generate potential fixes for each bug individually. The model processes each prompt, leveraging the historical and contextual information provided within the prompt, and outputs ten potential bug fixes. This output typically consists of function-level code, which is essential for the next steps in the pipeline.
  3. 3.
    Parsing the Model-generated Function-level Code. The code generated by Code Llama is parsed to extract the specific function-level code snippet. This is because, even though we explicitly instruct the model to generate only the fixed function-level code snippet, it often generates additional or unnecessary text, such as code explanations or unrelated code. To address this, we conducted multiple inferences with our prompt, observed the general output patterns, and developed regular expressions in our implementation to reliably isolate the desired function-level code snippet.
  4. 4.
    Replacing the Original Fixed Code with the Model-Generated Code. We begin by using Git commands to track the fixed snapshot of the project (commit V4 in Figure 1). Next, we locate the fixed file using the file path and name and identify the fixed function code by its start and end line numbers within the buggy function. To ensure a rollback option after test evaluation, we create a temporary backup (File_Backup) in the same directory. We then replace the original fixed code with the model-generated function-level code, so that we can later verify if this fix passes the test cases in the corresponding commit.
  5. 5.
    Running Test Cases in a Docker Environment. After reintegrating the code into the project, the next step is to validate the effectiveness of the generated fix. We install all dependencies for each bug in each project within a Docker environment, which ensures isolation and reproducibility during testing. The bug’s test cases are then executed, providing a consistent platform for evaluating the correctness of the generated code. After this test evaluation, we restore the original state by deleting the current file and renaming the backup file ((File_Backup)) to its original file name. To determine whether a bug is successfully fixed, we consider it resolved if at least one of the n samples (where n=10) generated by nucleus sampling passes the test cases, demonstrating functional correctness.
  6. 6.
    Calculating Pass@k as Evaluation Metric. In line with a previous study Chen et al. (2021); Du et al. (2023); Yu et al. (2024); Li et al. (2024b), we evaluate the functional correctness of programs by executing test cases to calculate the pass@k as shown in Formula 1. This step measures the success rate of the model-generated code over k attempts. Specifically, we generate $n$ code samples per query ($n\geq k$), then count the number of correct programs $c$ that pass the test cases ($c\leq n$), and calculate the Pass@k.
[…]
    For example, if 3 out of 10 generated samples pass the test cases, $c=3$ and $n=10$, allowing us to compute pass@k for $k=1,3,5,10$. We compute pass@k for each bug and aggregate the results across all bugs to derive the overall performance of HAFix and baseline approaches.
    Pass@k is chosen because it reflects the likelihood of the bug being fixed within k attempts, aligning with realistic bug-fixing scenarios where multiple solutions can be attempted. Compared to metrics used by previous works such as the number of bugs fixed or exact match (Jiang et al., 2023; Lu et al., 2021; Wang et al., 2021; Xia and Zhang, 2024), pass@k considers the distribution of correct fixes across multiple attempts rather than evaluating bug fixes in a binary manner. This provides a more nuanced measure of model effectiveness.
  7. 7.
    Conducting Statistical Test. To assess statistical significance, we apply the Friedman test, followed by Wilcoxon signed-rank tests for post-hoc analysis, comparing pass@k distributions across different heuristics and prompt styles. The Friedman test is chosen for its ability to detect differences across multiple related groups, while the Wilcoxon signed-rank test is well-suited for pairwise comparisons in non-parametric data. To complement statistical significance, we calculate the effect size using the Rank-Biserial Correlation, which quantifies the magnitude of pairwise differences. The Rank-Biserial Correlation ranges from -1 to 1, where values closer to -1 or 1 indicate stronger effects. These tests are consistently applied across RQ1, RQ2, and RQ3 to ensure uniform and robust statistical evaluation.
4.6 Inference Infrastructure and Hyper-parameters
For the infrastructure supporting our experiments, we selected the Nvidia A100 GPU with 80 GB of memory. It provides ample space to load our subject model as described in Subsection 4.3, which has a size of 12.55GB, along with the additional memory needed for processing large batches of data. Following prior works (Roziere et al., 2023; Li et al., 2024b), we employed nucleus sampling as the decoding strategy with a temperature of 0.4 and a top-p value of 0.95. Since hyper-parameter tuning was not our primary focus, we adopted these values from prior studies and left further exploration for future work. While the model generally produced stable outputs, it occasionally included unrelated text, such as code explanations or irrelevant snippets. To address this, we applied the extraction rules described in Subsection 4.5. These rules, developed by analyzing common output patterns, ensured the reliable isolation of function-level code snippets and maintained the stability of the evaluated fixed code samples.
5 Research Questions
5.1 RQ1: Do History-Augmented LLMs Improve Bug Fixing Compared to Models Without Historical Context?
5.1.1 Motivation
Large Language Models (LLMs) have demonstrated remarkable capabilities in SE tasks such as code generation (Lu et al., 2021; Chen et al., 2021; Du et al., 2023; Yu et al., 2024; Li et al., 2024b; Zhang et al., 2024a) and bug-fixing (Jiang et al., 2023; Zhang et al., 2024b; Fan et al., 2023; Xia et al., 2023; Li et al., 2024a). However, it remains uncertain whether incorporating historical context from software repositories, such as the blame commit of a bug, can further enhance their effectiveness in bug-fixing. The blame commit identifies the last modification to the buggy code, offering critical context for understanding its root cause. This context has been used in the MSR community for decades as a heuristic to identify the bug-introducing commit (Śliwerski et al., 2005; Hassan, 2006). As discussed in Section 3, examining the changes in the blame commit has long been used to understand the evolution of software bugs. This research question explores whether leveraging various history heuristics derived from the blame commit can improve LLMs’ bug-fixing performance.
5.1.2 Approach
To evaluate the impact of history heuristics, we use a baseline prompt inspired by GitHub Copilot’s prompt data (Copilot, 2024a), given that the latter is one of the most widely adopted coding assistant tools. The detailed prompt design for baseline is presented in Subsection 4.2.1. As discussed in Subsection 3.1, we empirically evaluate seven historical information into several heuristics: co-evolved functions’ names in the modified buggy files (CFN-modified), co-evolved functions’ names in all modified files (CFN-all), all functions’ names in the modified buggy file (FN-modified), all functions’ names in all modified files (FN-all), co-evolved files’ names (FLN-all), function code pairs (FN-pair), and file diff patches (FL-diff). When testing different heuristics, we always provide the baseline information first, then append the heuristic data to ensure a fair comparison of their impact on top of the baseline.
Additionally, we propose an aggregated approach named HAFix-Agg, as described in Subsection 3.2, which combines the results of all heuristics to assess its potential for improvement over the performance of individual heuristics. Note that in this research question, we used the Instruction prompt style for all experiments to maintain consistency.
5.1.3 Results
Figure 5: Pass@k comparison of individual heuristics and baseline for bug fixing performance using nucleus sampling (n=10).
Table 2: Pass@k (k=1, 5 and 10) for baseline and various history heuristics using nucleus sampling. The values higher than the corresponding baseline are shown in bold.
HAFix heuristics Pass@1 Pass@5 Pass@10
Baseline 19.41% 33.44% 39.22%
CFN-modified 17.65% 31.17% 37.25%
CFN-all 19.02% 34.41% 41.18%
FN-modified 13.53% 22.83% 25.49%
FN-all 14.12% 26.24% 33.33%
FLN-all 19.22% 35.55% 43.14%
FN-pair 15.29% 23.15% 25.49%
FL-diff 12.75% 22.17% 25.49%
FLN-all significantly improves bug-fixing performance compared to the baseline. Figure 5 illustrates the trend in Pass@k rates for the baseline and the seven heuristics across a range of k values from 1 to 10, providing a comprehensive view of their comparative performance on bug fixing. Notably, certain heuristics, particularly CFN-all and FLN-all, show consistent improvements over the baseline as k increases. Table 2 summarizes specific Pass@k results (k=1, 5, and 10, n=10) for the baseline and different heuristics using nucleus sampling. We selected these k values because they are commonly reported and provide a balanced perspective on performance at lower, mid, and higher thresholds (Du et al., 2023; Yu et al., 2024). Compared to the baseline’s rates of 33.44% at Pass@5 and 39.22% at Pass@10, CFN-all shows an improvement of approximately 3% at Pass@5 and 5% at Pass@10, while FLN-all improves by about 6% at Pass@5 and 10% at Pass@10 over the baseline.
(a) CFN-modified fixes 3 more bugs compared to the baseline.
(b) CFN-all fixes 6 more bugs compared to the baseline.
(c) FN-modified fixes 1 more bug compared to the baseline.
(d) FN-all fixes 4 more bugs compared to the baseline.
(e) FLN-all fixes 5 more bugs compared to the baseline.
(f) FN-pair fixes 2 more bugs compared to the baseline.
(g) FL-diff fixes 3 more bugs compared to the baseline.
Figure 6: Venn diagrams comparing the number of bugs fixed by the baseline (red) and the seven individual HAFix heuristics (green), with the overlapping region (brown) indicating bugs fixed by both the baseline and the heuristic. Numbers and percentages within each region denote the count and proportion of bugs fixed.
Figure 7: Venn diagram comparing the number of bugs fixed by the baseline (brown) and HAFix-Agg (green). The overlapping region represents bugs fixed by both the baseline and HAFix-Agg.
To assess whether CFN-all and FLN-all perform significantly better than the baseline, as well as to compare the performance of other heuristics, we conducted a Friedman test across the eight data groups of pass@k, with each group covering a range of k values from 1 to 10. The resulting p-value of 3.635e-12 indicates at least one significant difference among the groups. We then performed pairwise Wilcoxon signed-rank tests as post-hoc analysis to compare the baseline with CFN-all and FLN-all, applying a Bonferroni-corrected significance threshold of 0.0071 (0.05/7) for multiple comparisons. The comparison between CFN-all and the baseline yielded a p-value of 0.041, which does not meet the corrected threshold, while the comparison between FLN-all and the baseline resulted in a p-value of 0.006, confirming a statistically significant difference. To quantify the magnitude of these improvements, we calculated the effect size using the Rank-Biserial Correlation for the significant pairwise comparisons. The effect size for the comparison between FLN-all and the baseline was -0.51, indicating a moderate to strong improvement of FLN-all over the baseline. Additionally, other heuristics did not outperform the baseline in terms of Pass@k.
CFN-all and FLN-all fix the most unique bugs compared to the baseline. While the statistical analysis confirms a significant improvement for FLN-all, a closer examination reveals additional noteworthy trends. Figure 6 presents a series of Venn diagrams comparing the number of bugs fixed by the baseline and the seven history heuristics. Each Venn diagram consists of three overlapping regions: the red area represents the bugs fixed exclusively by the baseline, the green area represents the bugs fixed exclusively by the heuristic under evaluation, and the brown area in the middle represents the bugs fixed by both the baseline and the heuristic. A bug is considered fixed if at least one of the generated samples passes the test cases, as described in Subsection 4.5. As shown in the figure, CFN-all and FLN-all fix the most unique bugs compared to the baseline, addressing 6 and 5 additional bugs, respectively. FN-all and FL-diff follow, fixing 4 and 3 more bugs compared to the baseline, while CFN-modified also fixes 3 additional bugs compared to the baseline.
This leads us to consider an aggregated variant of HAFix (HAFix-Agg), which integrates the results of different history heuristics to enhance bug-fixing performance. As shown in the Venn diagram in Figure 7, HAFix-Agg not only fixes all 20 bugs addressed by the baseline but also resolves 9 additional bugs, highlighting a 45% improvement over the baseline. Table 3 further illustrates the number of bugs fixed by the baseline, HAFix heuristics, and HAFix-Agg. Out of the 51 bugs in our dataset, the baseline fixes 20, FLN-all fixes 22, while HAFix-Agg resolves 29, fixing 68.97% of the bugs compared to 39.22% for the baseline. While some heuristics show limited improvement individually, the combined effect of results from all heuristics offers a significant advantage over the baseline. This underscores the importance of incorporating diverse historical contexts, as they provide the model with a richer understanding and broader perspective for addressing bugs.
Table 3: Number and Percentage of bugs fixed by baseline and individual heuristics of HAFix. Bugs# represent the number of bugs being fixed.
HAFix Bugs# Percentage
Baseline 20 39.22%
CFN-modified 19 37.25%
CFN-all 21 41.18%
FN-modified 13 25.49%
FN-all 17 33.33%
FLN-all 22 43.14%
FN-pair 13 25.49%
FL-diff 13 25.49%
HAFix-Agg 29 56.86%
To assess HAFix-Agg’s performance relative to the baseline and the top individual heuristic (FLN-all) in terms of Pass@k, we conducted a dedicated experiment. HAFix-Agg’s results were obtained from seven heuristics together, with each heuristic generating 10 samples per bug via nucleus sampling, resulting in a total of 70 samples per bug. To ensure a fair comparison, the baseline and FLN-all were also run seven times respectively, generating a total of 70 samples per bug by executing their respective configurations seven times. This setup provided a consistent basis for calculating Pass@k across an expanded range of k values.
Figure 8: Pass@k comparison of HAFix-Agg, Baseline, and FLN-all using nucleus sampling (n=70).
HAFix-Agg consistently outperforms both the baseline and FLN-all when evaluated using nucleus sampling with n=70. Figure 8 illustrates the Pass@k trends for the baseline, FLN-all, and HAFix-Agg across k values, with k increasing in steps of 1. To assess statistical differences among these methods, we first applied the Friedman test, which revealed at least one significant difference between the three groups with a p-value lower than 2.2e-16. Following this, we conducted the Wilcoxon signed-rank test as a post-hoc analysis to identify pairwise differences. The p-values for the comparisons between HAFix-Agg and the baseline, as well as HAFix-Agg and FLN-all, were 2.9e-11 and 2.5e-06, respectively. After applying the Bonferroni correction, setting the significance threshold at 0.017 (0.05/3), both comparisons indicated statistically significant improvements for HAFix-Agg over the other methods. To quantify these improvements, we calculated the effect size using the Rank-Biserial Correlation. HAFix-Agg showed a strong improvement over the baseline (effect size = -0.46) and a moderate improvement over FLN-all (effect size = 0.33). As shown in the figure, all methods show an upward trend with increasing k, but HAFix-Agg consistently outperforms both the baseline and FLN-all, fixing additional bugs for higher k values, while the other approaches plateau much earlier. Notably, HAFix-Agg exceeds a 56% Pass@k rate as k increases, while the baseline plateaus around 45%. These findings highlight the substantial advantage of HAFix-Agg over both the baseline and the individual heuristics, emphasizing its effectiveness across the evaluated range of k values.
Handling Out-of-Memory (OOM) cases. During model inference, some heuristics encountered out-of-memory (OOM) issues due to larger prompt sizes or resource constraints. In total, 27 out of 408 cases (51 × 8) faced OOM issues. Bugs resulting in OOM errors were marked as failed in the Pass@k calculations, as no valid results were generated. This adjustment helps mitigate the limitation while maintaining a fair evaluation of HAFix.
Summary for RQ1 and other findings: 1. The FLN-all heuristic significantly improves bug-fixing performance, achieving a 10% (at Pass@10) higher improvement at fixing bugs compared to the baseline. 2. HAFix-Agg fixes 45% more bugs than the baseline, while also successfully addressing all the bugs fixed by the baseline. 3. With a larger number of samples, HAFix-Agg consistently outperformed both the baseline and FLN-all, demonstrating significant improvements in bug-fixing performance.
5.2 RQ2: How Do Different Prompt Styles Impact the Bug-Fixing Performance of History-Augmented LLMs?
5.2.1 Motivation
The effectiveness of LLMs can vary significantly depending on the structure and presentation of prompts (Xia and Zhang, 2022). This research question aims to investigate how different prompt styles influence the bug-fixing performance of the individual and aggregate HAFix variant HAFix-Agg. The primary motivation is to explore the potential for optimizing LLM performance by refining prompt structures to enhance bug-fixing outcomes. We aim to investigate how varying prompt styles affect the performance specifically for FLN-all and HAFix-Agg, the most effective heuristics identified in RQ1. By analyzing the impact of these styles, we seek to identify the most effective approach for fixing a greater number of bugs.
5.2.2 Approach
We utilize our baseline alongside the most promising bug-fixing approaches identified in RQ1, FLN-all, and HAFix-Agg, combined with three prompt styles: Instruction, InstructionLabel, and InstructionMask. These prompt styles vary in their levels of specificity and context structure, providing the potential to enhance the bug-fixing performance of our approach. The Instruction prompt provides explicit instructions on how to fix the bug, potentially improving clarity and precision in generating bug-specific fixed code. On the other hand, the InstructionLabel prompt combines instructions with the label of the buggy line of code. Finally, the InstructionMask prompt asks the model to complete missing segments of the buggy line of code that we have manually masked.
Since we already obtained the Instruction prompt results in RQ1, here we experimented with the InstructionLabel and InstructionMask prompts using nucleus sampling to identify the most effective prompt style for the baseline, FLN-all, and HAFix-Agg approaches, then compared to the RQ1 Instruction prompt results. For HAFix-Agg, we aggregated the 70 samples generated from the seven individual history heuristics (10 samples per heuristic, with each sample representing a single inference result from the LLM). Finally, we determined the best-performing prompt style for each approach and compared them with the baseline, FLN-all, and HAFix-Agg to identify the most effective prompt style.
To analyze the impact of different prompt styles on the baseline, FLN-all, and HAFix-Agg, we refer to the combination of an approach (baseline, FLN-all, or HAFix-Agg) with a specific prompt style (Instruction, InstructionLabel, or InstructionMask) as a configuration. Using the Pass@k metric, we evaluated and compared the performance of these configurations. Below, we present the results for each configuration, beginning with the baseline.
5.2.3 Results
Figure 9: Pass@k comparison of Baseline across different prompt styles using nucleus sampling (n=10).
Table 4: Pairwise comparisons of baseline across different prompt styles using Wilcoxon signed rank test. The p-value threshold is 0.017 (0.05/3) considering Bonferroni correction.
Baseline_InstructionMask Baseline_Instruction
Baseline_Instruction 0.002 -
Baseline_InstructionLabel 0.002 0.076
For the baseline heuristic, both Instruction and InstructionLabel significantly outperform InstructionMask. Figure 9 presents the Pass@k performance of the baseline across three prompt styles: Instruction, InstructionLabel, and InstructionMask. Across all evaluated k values, both Instruction and InstructionLabel demonstrate comparable Pass@k distributions, significantly outperforming InstructionMask. For instance, at k=10, the baseline with Instruction and InstructionLabel achieves a Pass@10 rate of around 39.22%. In contrast, InstructionMask lags behind with a Pass@10 rate of 13.73%.
To evaluate the statistical significance of these differences, a Friedman test was conducted, yielding a p-value of 0.00013. This result confirms the presence of at least one significant difference among the three prompt styles. Subsequently, post-hoc pairwise comparisons were performed using the Wilcoxon signed-rank test, with Bonferroni correction applied to account for multiple comparisons (0.05/3=0.017). Table 9 introduces the pairwise comparison results. The results show that both Instruction and InstructionLabel significantly outperform InstructionMask (p=0.002 for both comparisons), with a strong effect size of 0.55. However, the difference between Instruction and InstructionLabel is not statistically significant $(p=0.076>0.017)$. These findings suggest that both Instruction and InstructionLabel provide sufficient context to enhance baseline performance, whereas InstructionMask, which focuses on masked segment completion, is less effective in achieving high bug-fixing performance.
Figure 10: Pass@k comparison of FLN-all across different prompt styles using nucleus sampling (n=10).
Table 5: Pairwise comparisons of FLN-all across different prompt styles using Wilcoxon signed rank test. The p-value threshold is 0.017 (0.05/3) considering Bonferroni correction.
FLN-all_InstructionMask FLN-all_Instruction
FLN-all_Instruction 0.002 -
FLN-all_InstructionLabel 0.002 0.002
For FLN-all, Instruction is the best-performing prompt style achieving the highest Pass@k rates and significantly outperforming both InstructionLabel and InstructionMask. As FLN-all was identified in RQ1 as the best-performing individual heuristic, we extended the analysis to evaluate the impact of prompt styles on this heuristic. Figure 10 presents the Pass@k performance of FLN-all across the same three prompt styles. Similar to the baseline, FLN-all in both Instruction and InstructionLabel exhibit significantly higher Pass@k rates compared to InstructionMask. For instance, at k=10, FLN-all with Instruction and InstructionLabel achieves Pass@10 rates of 43.14% and 33.33%, respectively, while InstructionMask lags behind at 7.84%.
To evaluate the statistical significance of these differences, a Friedman test was performed, yielding a p-value of 4.54e-05, indicating at least one significant difference among the three prompt styles. Subsequently, post-hoc pairwise comparisons were performed using the Wilcoxon signed-rank test, with Bonferroni correction applied to account for multiple comparisons (0.05/3=0.017). Table 10 introduces the pairwise comparison results. The results show that FLN-all with Instruction significantly outperforms both FLN-all in InstructionLabel and InstructionMask (p=0.002 for both comparisons), with a strong effect size of 0.55 for each. Additionally, FLN-all in InstructionLabel significantly outperforms FLN-all in InstructionMask (p=0.002), also with a strong effect size of 0.55.
Figure 11: Pass@k comparison of HAFix-Agg across different prompt styles using nucleus sampling (n=70).
Table 6: Pairwise comparisons of HAFix-Agg across different prompt styles using Wilcoxon signed rank test. The p-value threshold is 0.017 (0.05/3) considering Bonferroni correction.
HAFix-Agg_InstructionMask HAFix-Agg_Instruction
HAFix-Agg_Instruction 3.6e-13 -
HAFix-Agg_InstructionLabel 3.6e-13 3.8e-13
HAFix-Agg with Instruction significantly outperforms both InstructionLabel and InstructionMask. After analyzing the individual heuristics (baseline and FLN-all), we proceeded to evaluate HAFix-Agg, which was confirmed in RQ1 to significantly outperform both baseline and FLN-all in the Instruction prompt style. Figure 11 shows how the Pass@k performance of HAFix-Agg across the three prompt styles: Instruction, InstructionLabel, and InstructionMask. Notably, Instruction achieves the highest Pass@k rates across all evaluated k values, outperforming both InstructionLabel and InstructionMask. For example, at k=10, HAFix-Agg with Instruction achieves a Pass@10 rate of 40.41%, while InstructionLabel reaches 31.67%, and InstructionMask remains around 11.55%.
To confirm the statistical significance of these differences, we conducted a Friedman test, which returned a p-value lower than 2.2e-16, indicating significant differences among them. Post-hoc Wilcoxon signed-rank tests with Bonferroni correction (0.05/3 = 0.017) further validated the results. Table 11 shows the pairwise comparisons, where HAFix-Agg with Instruction significantly outperforms both InstructionLabel and InstructionMask (p = 3.6e-13 and 3.8e-13), with a strong effect size of 0.51 for each. These results highlight that the clarity and explicitness of the Instruction prompt enable the model to leverage historical heuristics more effectively.
Figure 12: Pass@k comparison of best-performing baseline, FLN-all, and HAFix-Agg across different prompt styles using nucleus sampling (n=70).
Table 7: Pairwise comparisons of best-performing baseline, FLN-all, and HAFix-Agg using Wilcoxon signed rank test. The p-value threshold is 0.008 (0.05/6) considering Bonferroni correction.
Baseline_Instruction Baseline_InstructionLabel FLN-all_Instruction
Baseline_InstructionLabel 0.0098 - -
FLN-all_Instruction 3.7e-13 3.8e-13 -
HAFix-Agg_Instruction 2.9e-11 9.0e-13 2.5e-06
When comparing the best prompt styles of each approach, HAFix-Agg (Instruction) emerges as the overall best-performing configuration in the case of n=70, significantly outperforming both FLN-all and baseline approaches. Finally, we conducted a comprehensive comparison of the best-performing configurations from the previous analyses to determine the overall best approach. Figure 12 compares the Pass@k performance of the best-performing baseline (Instruction and InstructionLabel), FLN-all (Instruction), and HAFix-Agg (Instruction). Across all evaluated k values, HAFix-Agg consistently achieves the highest Pass@k rates, followed by FLN-all, with the baseline performing comparatively lower. For instance, at k=70, HAFix-Agg achieves a Pass@70 rate of 56.86%, while FLN-all reaches 50.98%. Both baseline configurations (Instruction and InstructionLabel) trail behind with Pass@10 rates to 47.06% and 49.02%, respectively.
A Friedman test confirmed significant differences among the configurations $(p<2.2e-16)$. Post-hoc Wilcoxon signed-rank tests with Bonferroni correction (0.05/6 = 0.008) further validated the results. Table 12 shows the pairwise comparisons, revealing that HAFix-Agg (Instruction) significantly outperforms all other configurations, including FLN-all (Instruction, p = 2.5e-06, effect size = 0.33) and the baseline (Instruction, p = 2.9e-11, effect size = 0.46; InstructionLabel, p = 9.0e-13, effect size = 0.50). FLN-all (Instruction) also surpasses both baseline configurations (Instruction, p = 3.7e-13; InstructionLabel, p = 3.8e-13; both effect size = -0.51). However, the difference between the two baseline configurations is not statistically significant $(p=0.0098>0.008)$. These results underscore the value of leveraging history-augmented heuristics alongside explicit instructions for maximizing bug-fixing performance.
Table 8: Number of out-of-memory (OOM) cases for HAFix across three prompt styles. Each style involves 408 (51 × 8) inference cases. The final row shows the total OOM cases and percentages for each style.
Prompt Styles Instruction InstructionLabel InstructionMask
Baseline 0 0 2
CFN-modified 0 0 4
CFN-all 2 2 7
FN-modified 1 1 4
FN-all 6 6 11
FLN-all 1 1 5
FN-pair 7 7 8
FL-diff 10 8 18
Percentage 27 (6.6%) 25 (6.1%) 59 (14.5%)
Handling Out-of-Memory (OOM) cases. Similar to RQ1, bugs resulting in OOM errors were labeled as failed for the Pass@k calculations to ensure a fair comparison across prompt styles. As shown in Table 8, InstructionMask had the highest rate of OOM cases (14.5%), followed by Instruction (6.6%) and InstructionLabel (6.1%), out of the total 408 inference cases for each prompt style. InstructionMask experienced the highest rate of OOM errors, particularly for heuristics like FL-diff and FN-all.
Summary for RQ2 and Key Findings: 1. The Instruction prompt style significantly outperforms InstructionLabel and InstructionMask for HAFix-Agg. 2. FLN-all with Instruction consistently outperforms the baseline across all evaluated prompt styles, highlighting the combined effectiveness of incorporating history through FLN-all and the clarity provided by the Instruction prompt style. 3. The Instruction prompt style achieves the highest performance overall, significantly outperforming other prompt styles across various approaches and configurations.
5.3 RQ3: What Is the Cost and Efficiency of History-Augmented LLMs on Bug Fixing?
5.3.1 Motivation
While HAFix-Agg in Instruction prompt style significantly improves bug fixing, it is essential to examine the trade-offs between performance, cost, and time efficiency. Unlike single-heuristic approaches, HAFix-Agg requires seven separate inference requests, one for each heuristic, resulting in a sevenfold increase in inference time and financial cost. This raises important questions about the practicality of deploying HAFix-Agg at scale. Specifically, RQ3 focuses on inference time and inference price. Inference time measures the model’s time efficiency, indicating how quickly it can generate results. Inference price, on the other hand, represents the price required to process input and output tokens during bug fixes, an important factor when considering large-scale deployment.
5.3.2 Approach
First, we use the inference time across different bugs and heuristics which measures the duration from sending the prompt query to receiving 10 sample results using nucleus sampling. To ensure consistency, all experiments are conducted using identical infrastructure, including the same GPU environment and network, during a continuous, uninterrupted computation period. This minimizes external variables and maintains experimental stability.
Second, we adopt the inference price metric that estimates the monetary cost of bug-fixing inference by referencing the pricing structure of a popular commercial model GPT-3.5-turbo-instruct (OpenAI, 2024). Although we utilize an open-source, locally hosted model for our experiments, this approach allows us to approximate real-world deployment costs more accurately. This cost is calculated based on the number of input and output prompt tokens for a single inference (covering 10 output samples by nucleus sampling). For nucleus sampling (n=10 in our experiment), the cost formula for fixing a single bug is provided in Equation 2.
[…]
Where:
  • $T_{\text{input}}$ is the number of input tokens.
  • $P_{\text{input}}$ is the price per input token.
  • $T_{\text{output}}(i)$ denotes the number of output tokens for the $i$-th sample.
  • $P_{\text{output}}$ is the price per output token.
HAFix-Agg aggregates the inference results from different individual heuristics. Performing one inference of HAFix-Agg involves executing multiple heuristics, with the total cost depending on the sequence and the point at which a bug is successfully fixed during the heuristic execution process. We define four scenarios to calculate cost and time efficiency. The key differences among these scenarios relate to when bug fixing stops within the execution sequence of heuristics and the order in which the heuristics are executed. The four scenarios are defined as follows:
  • Exhaustive: Executes baseline and all heuristic methods for every bug, regardless of order. This means that all seven individual heuristics of HAFix yield a separate inference request, each resulting in 10 sample bug fixes.
  • ES (EarlyStop): Goes through the baseline and individual heuristics in the order shown in Table 2, and stops as soon as one good fix (i.e., passing the test case) has been generated. The execution sequence is: Baseline, CFN-modified, CFN-all, FN-modified, FN-all, FLN-all, FN-pair, FL-diff.
  • ES-AccSorted: Based on the idea of ES, the order of baseline and heuristics is determined based on the number of bugs fixed across our dataset, as shown in Table 3. Heuristics that fix more bugs earlier are prioritized. The execution sequence is: FLN-all, CFN-all, Baseline, CFN-modified, FN-all, FN-modified, FN-pair, FL-diff.
  • ES-UniSorted: Also based on the idea of ES, but prioritizes the heuristics based on the number of bugs uniquely fixed by each heuristic compared to the baseline, as shown in Figure 6. The execution sequence is: Baseline, CFN-all, FLN-all, FN-all, CFN-modified, FL-diff, FN-pair, FN-modified.
We apply these four scenarios to calculate the inference price and time efficiency, aiming to identify any trade-offs between performance, price, and inference time across the different scenarios. This RQ focuses exclusively on HAFix in the Instruction prompt style, as it demonstrated the best performance. For cases encountering out-of-memory issues (27 cases, representing 6.6% of the total 51 × 8 cases, as shown in Table 8) and failing to generate fixed code, their inference times are excluded from the analysis to avoid potential bias, as these cases typically require significantly more time. However, their inference prices are included by accounting only for the cost of input tokens, with the cost of output tokens set to zero due to the absence of generated outputs. This helps mitigate the limitation while maintaining a fair evaluation.
5.3.3 Results
Figure 13: Distribution of inference time of bugs across different cost scenarios.
Table 9: Pairwise comparisons of inference time using Wilcoxon signed rank test for different cost scenarios. The p-value threshold is 0.008 (0.05/6) considering Bonferroni correction.
Pairwise Compare Exhaustive ES ES-AccSorted
ES 4.0e-06 - -
ES-AccSorted 4.0e-06 0.737 -
ES-UniSorted 2.7e-06 0.076 0.387
The Exhaustive scenario is significantly more time-consuming than all other scenarios. Figure 13 presents the distributions of inference time across all bugs for the four cost scenarios: Exhaustive, ES, ES-AccSorted, and ES-UniSorted. The Exhaustive scenario has the highest median inference time of 303 seconds (s) and a broader range, reflecting longer and more variable processing times. In contrast, the early stopping scenarios, including ES (145s), ES-AccSorted (104s), and ES-UniSorted (145s), exhibit much lower and more consistent inference times. Additionally, the difference between the sorted (ES-AccSorted and ES-UniSorted) and unsorted (ES) approaches is minimal.
Figure 14: Distribution of inference time for each bug across different heuristics in the Exhaustive scenario.
To explore these differences further, we performed a Friedman test, which confirmed a statistically significant variance in inference time across these scenarios (p-value = 2.148e-12). A subsequent post hoc analysis using the pairwise Wilcoxon test in Table 9 reveals that the Exhaustive scenario is significantly more time-consuming than all other scenarios, with p-values of 4.0e-06 (effect size = -0.20) or 2.7e-06 (effect size = -0.18), both of which are below the Bonferroni-corrected threshold of 0.008 (0.05/6). These results highlight that early stopping significantly reduces inference time compared to the Exhaustive scenario, with minimal differences between sorted and unsorted scenarios.
FN-pair is the most time-efficient heuristic. To gain deeper insights into the impact of individual heuristics on inference time, we focus on the Exhaustive scenario, as it involves executing all individual heuristics. In this scenario, each bug processed by each heuristic is associated with a specific inference time value. This allows us to identify the most time-efficient heuristics and those with higher time demands. Figure 14 shows how the distribution of inference times for the various heuristics in the Exhaustive scenario. The median inference times range from 31 seconds for FN-pair, the most time-efficient heuristic, to 44 seconds for FN-all, the slowest heuristic. FLN-all, the best-performing individual heuristic identified in RQ1, achieves a median inference time of 37.5 seconds, striking a balance between time efficiency and performance.
Table 10: Pairwise comparisons of inference time using Wilcoxon signed rank test for different heuristics. The p-value threshold is 0.0018 (0.05/28) considering Bonferroni correction.
Baseline CFN-modified CFN-all FN-modified FN-all FLN-all FN-pair
CFN-modified 0.0413 - - - - - -
CFN-all 0.0004 0.0570 - - - - -
FN-modified 1.7e-05 0.0368 0.3324 - - - -
FN-all 0.0039 0.0324 0.1274 0.2859 - - -
FLN-all 0.0153 0.2457 0.1331 0.1217 0.1242 - -
FN-pair 0.6758 0.0829 0.0025 0.0006 0.0003 0.0349 -
FL-diff 0.4878 0.9365 0.7007 0.4769 0.4720 0.9885 0.1167
To determine whether these differences are statistically significant, we applied the Friedman test, which indicated significant variance in inference times among the heuristics (p-value = 4.736e-07). Pairwise comparisons using the Wilcoxon test, as shown in Table 10, reveal that FN-pair has significantly lower inference times compared to FN-modified (effect size = -0.33) and FN-all (effect size = -0.34). However, FN-pair does not differ significantly from the other heuristics. These findings demonstrate that FN-pair is the most time-efficient heuristic, while FLN-all offers a well-balanced trade-off between time efficiency and performance.
Figure 15: Percentage of bugs being fixed (Table 3) vs. total inference time across different heuristics in the Exhaustive scenario. The inference time reflects the total time required to process all bugs for each heuristic.
Heuristics achieving higher bug-fixing percentages incur longer inference time. To further explore the relationship between heuristic performance and inference time efficiency, Figure 15 illustrates the percentage of bugs fixed in relation to inference time for different heuristics in the Exhaustive scenario. The y-axis represents the percentage of bugs being fixed, while the x-axis denotes inference time in seconds. The figure reveals a clear trade-off: heuristics such as FLN-all, CFN-all, and Baseline fix a higher percentage of bugs but require similar longer inference times, reflecting their computational intensity. In contrast, FN-pair and FL-diff exhibit shorter inference times but fix a lower percentage of bugs, indicating reduced effectiveness.
Figure 16: Distribution of inference price of bugs across different cost scenarios.
Table 11: Pairwise comparisons of inference price using Wilcoxon signed rank test for different cost scenarios. The p-value threshold is 0.008 (0.05/6) considering Bonferroni correction.
Pairwise Compare Exhaustive ES ES-AccSorted
ES 4.0e-06 - -
ES-AccSorted 4.0e-06 0.829 -
ES-UniSorted 2.7e-06 0.058 0.299
Early stopping strategies have a substantial impact on cost reduction. Figure 16 presents the distribution of inference prices for bugs across the four scenarios: Exhaustive, ES, ES-AccSorted, and ES-UniSorted. The median inference prices per bug are \$0.07 for the Exhaustive scenario, \$0.03 for ES, \$0.02 for ES-AccSorted, and \$0.03 for ES-UniSorted. These values highlight that the Exhaustive scenario is considerably more costly than the others, with early-stopping scenarios showing significantly lower median costs.
To statistically confirm these observations, we performed a Friedman test, which revealed a significant variance in inference price across the scenarios (p-value = 3.208e-12). A subsequent post hoc analysis using the pairwise Wilcoxon test identified statistically significant differences between the Exhaustive scenario and all other scenarios, as shown in Table 11. The pairwise p-values (4.0e-06 and 2.7e-06) are well below the Bonferroni-corrected threshold of 0.008 (0.05/6), with effect sizes of -0.20 and -0.18, indicating significant differences in the distributions. Referring to the median values, it is evident that the Exhaustive scenario incurs substantially higher costs (more than double) than the early stopping scenarios.
These results underscore the effectiveness of the early-stop strategy in reducing inference costs. By halting the process once a bug is fixed, early stopping avoids unnecessary computation, leading to lower median costs and reduced variability. The Exhaustive scenario incurs significantly higher inference price than all other cost scenarios, with early stopping showing a substantial impact on cost reduction.
Table 12: The total inference price of baseline and different heuristics across different cost scenarios.
Exhaustive ES ES-AccSorted ES-UniSorted
Baseline 0.53 0.53 0.30 0.53
CFN-modified 0.53 0.37 0.28 0.27
CFN-all 0.54 0.33 0.32 0.36
FN-modified 0.56 0.29 0.28 0.27
FN-all 0.53 0.27 0.27 0.27
FLN-all 0.53 0.26 0.53 0.28
FN-pair 0.41 0.19 0.19 0.17
FL-diff 0.54 0.30 0.30 0.30
Total 4.18 2.54 2.46 2.44
ES, ES-AccSorted, and ES-UniSorted have similar costs, suggesting that sorting strategies have a minimal impact on inference price. Having examined the cost distribution at the bug level, we next investigate the cost distribution at the heuristic level and the total cost differences among the four scenarios. Table 12 introduces the total inference price for each scenario, showing the contribution of both the baseline and various heuristics across the four scenarios. As shown in the table, the total inference prices for the early stopping scenarios, namely ES (\$2.54), ES-AccSorted (\$2.46), and ES-UniSorted (\$2.44), are all lower than that of the Exhaustive scenario (\$4.18). While Exhaustive is nearly twice as expensive, the absolute price difference is small, making it a viable option when cost is not a primary concern. However, as shown in Table 13, the total inference time across all bugs is substantially higher for Exhaustive, making it less practical for time-sensitive applications. Additionally, while early stopping significantly reduces costs, the additional cost savings from sorting strategies based on either bug-fixing performance or unique fixes are minimal, with only slight differences between ES, ES-AccSorted, and ES-UniSorted.
The Wilcoxon test results in Table 11 further support this observation. All p-values comparing ES, ES-AccSorted, and ES-UniSorted exceed the Bonferroni-corrected threshold of 0.008, indicating no statistically significant differences between these scenarios. Moreover, Table 12 highlights that certain heuristics, such as CFN-all and FN-modified, contribute significantly to the total cost in the Exhaustive scenario. However, these contributions are drastically reduced in the early stopping scenarios, emphasizing the effectiveness of early stopping as a cost-saving strategy.
Figure 17: Cost trends of different heuristics across four cost scenarios for HAFix-Agg. The positive y-axis represents the inference price for successfully fixed bugs, while the negative y-axis corresponds to the inference price for attempted but failed fixes. Numerical values above each bar indicate the number of bugs successfully fixed on top of the preceding heuristic. The red and blue lines depict the changing trends in the number of successfully fixed bugs and failed fixes, respectively. Bars are displayed in the order of heuristic execution within each cost scenario, from left to right.
Prioritizing heuristics on top of the early stop strategy applied is not highly effective. With the overall cost differences established, we now explore how the cost trend evolves during the bug-fixing process within each scenario. Figure 17 illustrates the cost distribution of various heuristics across the four cost scenarios for HAFix-Agg, showing the cost of successfully fixed bugs (above the x-axis) and unfixed bugs (below the x-axis). As seen in the figure, prioritizing heuristics such as Baseline and CFN-all fix the most bugs but didn’t show much difference among ES, ES-AccSorted, and ES-UniSorted scenarios. In contrast, when comparing Exhaustive with the other three scenarios, we found early stopping avoids running inefficient heuristics and reduces costs clearly. This is because, in the Exhaustive scenario, all heuristics are applied regardless of their time efficiency, leading to higher costs without a proportional increase in successful fixes. This trend further suggests that an early stopping strategy is useful, while on top of this, prioritizing the more effective ones doesn’t show a big difference.
Figure 18: Percentage of bugs being fixed (Table 3) vs. total inference price across different heuristics in the Exhaustive scenario.
FLN-all, CFN-all, and Baseline achieve higher bug-fixing performance but incur similar greater inference costs. After understanding the cost trends across heuristics, it is also crucial to explore how cost correlates with heuristic performance. Figure 18 depicts the relationship between the percentage of bugs fixed and inference price for various heuristics in the Exhaustive scenario. The figure reveals that heuristics such as FLN-all, CFN-all, and Baseline achieve higher bug-fixing performance but incur similar greater inference costs particularly compared to FN-pair. Conversely, heuristics like FN-pair exhibit lower performance while maintaining the lower price, reflecting reduced time efficiency in balancing performance and cost. This trade-off underscores the importance of strategically selecting heuristics to optimize both bug-fixing performance and cost, particularly in scenarios where computational resources are limited.
Summary for RQ3 and other findings: 1. Early stopping significantly reduces inference time compared to the Exhaustive scenario. The differences between sorted and unsorted approaches are minimal, indicating that sorting strategies provide little additional efficiency. 2. Early stopping significantly influences the speed at which bugs are fixed by HAFix-Agg. However, applying additional prioritization heuristics on top of early stopping does not demonstrate substantial effectiveness. 3. A trade-off exists between heuristics performance and inference resources, with higher-performing heuristics requiring greater inference time. Despite this, heuristics generally exhibit similar inference prices with minor variations in consistency.
6 Threats to Validity
6.1 Internal Validity
One potential threat to the internal validity of our study is the absence of LLM-based APR tools as a baseline for comparison. This decision stems from our primary focus on investigating the historical aspect of bug-fixing, specifically whether history data can enhance the performance of LLMs in bug-fixing tasks. For this reason, we designed a baseline inspired by GitHub Copilot, one of the most widely used coding assistants, to ensure relevance and practical applicability. We approach this study from the perspective of prompt engineering, with insights that may extend to other LLM applications in software engineering, such as code generation.
Another internal threat involves potential data leakage when training the Large Language Models. The commit messages and bug descriptions mined from GitHub might contain explicit hints or detailed information on how the bug was fixed, which could influence the LLM’s ability to generate correct repairs. To address this, we manually reviewed all mined texts and two people discussed and removed those that contained repair-specific information. While this manual filtering reduces the risk of leakage in the model inference stage, the data leakage during training may still exist, i.e., the LLM might have seen the issue report from collected open-source training datasets. However, we believe our manual work mitigates the issue in the context of our experiments during model inference.
A further threat relates to the out-of-memory (OOM) issues encountered during inference for some particularly large or complex bugs. These OOM errors could potentially be avoided by using more powerful hardware or distributing the task across multiple GPUs. Nevertheless, to maintain consistency across experiments and ensure a fair comparison, we conducted all experiments on a single A100 GPU with 80GB of memory. We believe this configuration provides a reasonable balance between computational capacity and accessibility, as it is capable of handling the majority of cases in practical scenarios.
Regarding model selection, we used the 7b instruct-tuned version from the Code Llama family, chosen for its balanced trade-off between performance and resource efficiency, with a size of 12.55GB. While the 13b and 34b variants may offer improved performance, they require significantly more computational resources and memory. This choice could limit the baseline’s performance, as larger models may generate higher-quality code, posing a potential threat to internal validity by not representing the optimal performance.
Lastly, in RQ3, we calculate the runtime and price of HAFix solely during the model inference stage, excluding the cost of running test cases. We observed that most of the time required for running test cases is consumed by installing the dependent libraries, averaging around 6 minutes per bug, while the actual execution of the test cases typically takes no more than a second, given that they usually consist of just a few lines of code. In practice, setting up the test environment is generally a one-time action, as developers typically maintain a stable development environment and do not reinstall libraries for each test. Regarding inference price, resources for model inference on GPUs are usually much more expensive than running test cases on a local CPU environment. Therefore, we believe that the cost and efficiency analyzed in RQ3 were underestimated but cover the main aspects of the real cost.
6.2 External Validity
One threat to external validity in our study lies in the choice of model. We focused our experiments exclusively on the Code Llama model, which, while popular and representative of open-source LLMs for code generation, may not generalize to other models, particularly more recent or proprietary systems. Different models may exhibit varying performance characteristics, such as handling different types of contexts or bugs more effectively. However, given the wide adoption and strong performance of Code Llama in the research community (Roziere et al., 2023), we consider it a suitable and relevant choice for evaluating our approach.
Additionally, we chose to evaluate our approach using the BugsInPy dataset, a relatively new and popular APR dataset, rather than the more established Defects4J dataset (Just et al., 2014). This decision aligns with our focus on Python bugs, which is the main scope of this paper. Additionally, Martinez et al. (2017) noted that the test suites in Defects4J are often weak and cannot guarantee the functionality completeness of the program under testing. Recent findings (Rafi et al., 2023) revealed that 77% of Defects4J’s fault-triggering tests contain developer knowledge added after bug reporting, introducing biases that may overestimate repair performance. The detailed reasons for choosing BugsInPy are detailed in Subsection 4.1. However, this choice, combined with the use of a dataset comprising 51 single-line bugs from 11 open-source projects, may limit the generalizability of our findings. We investigated the recent works and found Prenner et al. (2022) evaluate 40 bugs in both Java and Python, Kolak et al. (2022) evaluate 72 Python bugs and Chen et al. (2024) evaluate 124 bugs, most of recent existing works focus on single-line bugs (Xia and Zhang, 2022; Ye et al., 2022; Jiang et al., 2023; Prenner and Robbes, 2024). While our dataset is diverse, it may not fully represent the complexity of real-world bug-fixing scenarios, particularly those involving multi-line or multi-hunk bugs. Extending our study to other programming languages and larger, more varied datasets could yield additional insights and broaden the applicability of our approach. Future work should explore the scalability of history-augmented LLMs across different programming languages and more complex bug scenarios such as multi-line or multi-hunk bugs, to validate their effectiveness in diverse contexts.
6.3 Construct Validity
Regarding construct validity, a notable limitation stems from the inherent non-determinism in the outputs of LLMs. As LLMs generate results probabilistically, repeated runs might yield slightly different outputs, potentially leading to variations in pass@k results. To address this, we utilized nucleus sampling to generate 10 different code samples for each bug during inference. We also run the baseline, FLN-all 7 times for three prompt styles separately as detailed in RQ1 5.1.3 and RQ2 5.2.3, and conduct a stability analysis of one of the repeated runs in the following Discussion Section 7. These approaches help reduce the impact of randomness in the generation process and provide a more comprehensive evaluation of the model’s capabilities, although some variation in outcomes may still occur.
7 Discussion and Future Work
To assess the stability of the model’s inference results, we used the evaluation data from RQ1 (Subsection 5.1) which repeated running the baseline (in Instruction prompt style) 7 times under identical conditions. The coefficient of variation (CV), as summarized in Table 13, ranged from 4.37% to 8.03%, indicating low to moderate variability. This level of consistency is expected in stochastic processes like large language model (LLM) inference, confirming the robustness of the reported findings.
Table 13: Summary of the stability analysis for pass@k values across 7 repeated runs of the baseline in the instruction prompt style, including the mean pass@k and coefficient of variation (CV) for each k value.
k Mean Pass@k Coefficient of Variation
1 19.38% 4.37%
2 25.50% 6.10%
3 29.32% 6.59%
4 32.04% 6.85%
5 34.09% 7.09%
6 35.68% 7.31%
7 36.93% 7.49%
8 37.90% 7.66%
9 38.66% 7.83%
10 39.22% 8.03%
In our experiments, FLN-all, which involves using the co-evolved files’ names in the blame commit, showed the highest bug-fixing performance among all individual heuristics. This result highlights the value of co-evolution as an essential contextual feature. Co-evolved files often share dependencies or functional relationships, and their inclusion helps the model understand the broader operational context of the buggy code. Such interconnected nature in modern codebases, where changes in one file can propagate to others, makes holistic context indispensable for accurate bug fixes.
Moreover, the aggregated HAFix variant HAFix-Agg, builds on this insight by combining the results of multiple individual heuristics, ensuring the highest bug-fixing performance on top of the bugs fixed by the baseline. Lastly, the HAFix architecture, which organizes results from diverse contextual data components, demonstrates its potential as a modular, robust, and scalable strategy for future LLM-based, history-augmented bug fixing.
In terms of prompt styles, the Instruction prompt consistently outperformed other styles, showing significant improvements in bug-fixing performance. The likely reason is that this style provides the model with clear, direct instructions, which allows it to focus on generating bug-specific solutions without ambiguity. In practice, this finding implies that, when applying LLMs to bug-fixing tasks, structuring the prompt with explicit instructions should be the preferred approach for maximizing performance.
Additionally, despite the Instruction style success, there is always a trade-off between bug-fixing performance, inference price, and time efficiency. Models like Code Llama, while open-source and free, can still incur high computational costs, especially when generating multiple samples or handling large codebases. Based on our cost analyses, we found that scenarios such as ES-AccSorted and ES offer a better balance by reducing the inference price while maintaining competitive performance, while the price difference among them is not much. In practice, developers must carefully consider the trade-offs between the desire for maximum performance and the constraints of computational resources, such as GPU memory and inference time.
Moreover, the inference time data highlights an important consideration for real-world applications. While settings like FN-pair offer lower median inference times, this comes at the cost of lower performance. In contrast, settings like FLN-all deliver better performance but demand significantly longer inference times. This suggests that different settings should be prioritized based on the specific requirements of the task at hand: scenarios that demand higher performance may need to tolerate longer runtimes, while time-sensitive tasks could benefit from faster, albeit less accurate, configurations.
For future work, our approach is specialized for single-line bugs, potentially limiting its effectiveness when applied to more complex scenarios, such as multi-line or multi-hunk bugs. While current state-of-the-art (SOTA) LLM-based APR tools primarily focus on single-line bugs (Xia and Zhang, 2022; Ye et al., 2022; Jiang et al., 2023; Prenner and Robbes, 2024), future work will aim to extend the applicability of our approach to more complex cases, including multi-hunk bugs.
Moreover, exploring deeper historical information beyond blame commit information is possible. Although this work focused on utilizing co-evolved files’ names from the blame commit, other recent commit information may provide additional insights that can further improve bug-fixing performance. For example, exploring commit histories preceding or following the bug-introducing commit may reveal patterns or relevant code changes that can inform the fix. Future research could explore methods to extract and integrate this broader historical information into the bug-fixing process, potentially improving the model’s understanding of how bugs emerge and evolve.
Furthermore, although this study focused on bug-fixing, the history-based heuristics that we developed could be evaluated on other software engineering tasks, such as code generation or completion. For these tasks, the co-evolution of files and commit histories might also provide useful context, helping LLMs to generate more contextually aware and consistent code. Future work should investigate the effectiveness of these heuristics in different tasks and assess how well they generalize across various code-related challenges.
Last but not least, as LLMs are deployed at scale, managing the cost and efficiency of inference becomes increasingly important. Developing better cost models that account for both token usage and hardware limitations will be essential to ensure the practical scalability of LLM-based bug fixing.
8 Conclusion
Inspired by the foundations of mining software repositories, this study explores the integration of seven different ways (and one aggregated variant HAFix-Agg) of adding historical context into LLM-based bug fixing, evaluates the impact of different LLM prompt styles, and investigates trade-offs between bug-fixing performance, cost, and time of efficiency. The results reveal critical insights that advance the understanding of how LLMs can be optimized for practical use in bug fixing.
The incorporation of historical heuristics, particularly through FLN-all, demonstrates significant improvements in bug-fixing performance compared to a baseline inspired by GitHub Copilot. Our aggregated approach, HAFix-Agg, extends these improvements by leveraging the combined strengths of multiple heuristics, achieving a 45% increase in bugs fixed while addressing all bugs resolved by the non-history-based baseline. This finding highlights the importance of historical context as a valuable addition to LLM prompts for understanding and addressing bugs, reflecting the way in which historical software engineering data made great strides towards better software analytics early on in the mining software repositories domain.
The analysis of prompt styles underscores the critical role of prompt design in influencing LLM effectiveness. Among the three styles evaluated, Instruction consistently outperformed InstructionLabel and InstructionMask, demonstrating its ability to provide clarity and explicitness that enables the model to make optimal use of historical heuristics. This finding establishes Instruction as the preferred style for crafting prompts in history-augmented bug fixing.
In evaluating the trade-offs between performance, cost, and time efficiency, early stopping strategies emerged as a practical solution for reducing inference time and price without compromising effectiveness. Scenarios like ES-AccSorted and ES-UniSorted prioritized high-performing heuristics but their benefits over the basic ES strategy were marginal. In contrast, later-stage heuristics offered diminishing returns, contributing fewer fixes at higher costs and longer processing times.
These findings collectively provide actionable insights into optimizing LLM-based bug fixing. By integrating historical heuristics, employing effective prompt designs, and leveraging cost-efficient execution strategies, developers can enhance both the practicality and scalability of automated bug-fixing systems over using individual, history-unaware approaches. Future research could explore extending these methods to more complex types of bugs, additional programming languages, and other software engineering tasks such as code generation, to further validate and refine the role of historical software context in LLMs for software analytics tasks.
Data Availability Statement The experiment code and dataset constructed of this study are available in the replication package (Shi et al., 2024).
Conflict of Interest: All authors certify that they have no affiliations with or involvement in any organization or entity with any financial interest or non-financial interest in the subject matter or materials discussed in this manuscript.
Funding: This study was funded by NSERC.
Ethical Approval: This article does not contain any studies with human participants or animals performed by any of the authors.
Appendix A Prompt Example
<s>[INST] <<SYS>> You are a helpful and honest code assistant expert in fixing the buggy code in Python. I mined a buggy code snippet and its related information from GitHub. I will provide you with the project name, buggy file name, buggy function name, the date time, the current version of this buggy code snippet, the corresponding bug description that might indicate how this buggy code should be fixed, and the buggy line content that might suggest where this buggy code should be fixed. Please only generate the fixed code snippet of this buggy code, dont explain any other things. Please wrap your fixed code snippet between ‘‘‘python and ‘‘‘ <</SYS>> # The project name: luigi # The buggy file name: scheduler.py # The buggy function name: get_pending_tasks # The buggy code snippet: def get_pending_tasks(self, state): """ Get PENDING (and RUNNING) tasks for this worker. You have to pass in the state for optimization reasons. ␣␣␣␣""" if len(self.tasks) < state.num_pending_tasks(): return six.moves.filter(lambda task: task.status in [PENDING, RUNNING], self.tasks) else: return state.get_pending_tasks() # The bug description: Filters tasks in second branch of Worker.get_pending_tasks (#1849) When a worker has many DONE tasks, get_pending_tasks may switch to using state.get_pending_tasks in order to speed up the process. This can include pending tasks not owned by the worker, invalidating the result and causing functions like is_trivial_worker to return erroneous results. To fix this, we simply filter the results of state.get_pending_tasks to remove any tasks that dont include this worker. # The buggy line content: return state.get_pending_tasks() # The fixed code snippet: [/INST]
Listing 1: A real example of a baseline prompt for a bug from the Luigi project. [^1] https://github.com/spotify/luigi/commit/3c55acd2cd5cf9c6c760bec5bb3159e0bc48a614 The prompt is designed based on the template from the official Code Llama documentation. [^2] https://www.llama.com/docs/model-cards-and-prompt-formats/meta-code-llama/ The system prompt is enclosed within <<SYS>>, and the different components are structured by underscores.
Appendix B Representative Example of Dataset
"bug_id": {
"project_name": "luigi",
"project_url": "https://github.com/spotify/luigi.git",
"bugsinpy_id": "10",
"is_single_line": true,
"buggy_line_location": 305,
"buggy_line_content": " return state.get_pending_tasks()",
"in_function": true,
"commit": {
"commit_id": "3c55acd2cd5cf9c6c760bec5bb3159e0bc48a614",
[…]
"commit_parent": "f538d1b3d473d542a19d508e5f7e0809b1dfe5ef",
"commit_date": "2016-09-12 09:51:39",
[…]
"function": {
"function_name": "get_pending_tasks",
"function_parent": "Worker",
"function_before_start_line": 295,
"function_before_end_line": 305,
"function_after_start_line": 295,
"function_after_end_line": 305,
"function_before_token_count": 54,
"function_after_token_count": 72,
[…]
"file": {
"file_name": "scheduler.py",
"file_path": "luigi/scheduler.py",
"file_nloc": 952,
"file_complexity": 375,
"file_token_count": 7424
Listing 2: A representative example of a JSON-formatted dataset entry. This example shows metadata about a bug.